the electronic force between two-points charge with a magnitude of-800nC and is 15N.how far apart are the two charges from each other in centimeters? with solution.

1
Expert's answer
2022-03-09T12:29:35-0500

Force

"F=\frac{kq_1q_2}{r^2}"

"r^2=\frac{kq_1q_2}{F}"


"r^2=\frac{9\times10^9\times800\times10^{-9}\times800\times10^{-9}}{15}=3.84\times10^{-04} m"

"r=0.0195m"

"r=1.95cm"